OpenStudy (anonymous):

Suppose that a, b, and c are complex numbers such that |a|=|b|=|c|=1 and a+b+c=0. Prove that |a-b|=|b-c|=|c-a| and discuss the geometric meaning. The hint I received was to set a=1.....but I don't see where that is going.

OpenStudy (kirbykirby):

I am almost done typing.. I'm sorry this is kind of long!!

OpenStudy (kirbykirby):

This is a bit long.. Maybe someone else has a shorter solution Let: \(a=x_1+iy_1\) \(b=x_2+iy_2\) \(c=x_3+iy_3\) Then: \(|a|=\sqrt{x_1^2+y_1^2}=1\implies x_1^2+y_1^2=1\) \(|b|=\sqrt{x_2^2+y_2^2}=1\implies x_2^2+y_2^2=1\) \(|c|=\sqrt{x_3^2+y_3^2}=1 \implies x_3^2+y_3^2=1 \) and: \(a+b+c=0 \\ \implies (x_i+iy_1)+(x_2+iy_2)+(x_3+iy_3)=0\\ \implies (x_1+x_2+x_3)+i(y_1+y_2+y_3)=0 +i0\\ \\~ \\ \implies x_1+x_2+x_3=0 ~\text{ and }~y_1+y_2+y_3=0 \) Which implies: (I'll call these equations by\( (*)\) \(x_1=-(x_2+x_3)\\ x_2=-(x_1+x_3)\\ x_3=-(x_1+x_2)\) and the same idea for the \(y\)'s So: \[ \begin{align}|a-b|&=|(x_1+iy_1)-(x_2+iy_2)|\\&=|(x_1-x_2)+i(y_1-y_2)| \\&=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\\&=\sqrt{x_1^2-2x_1x_2 +x_2^2+y_1^2-2y_1y_2+y_2^2}\\&=\sqrt{\underbrace{x_1^2+y_1^2}_{=1} +\underbrace{x_2^2+y_2^2}_{=1}-2(x_1x_2+y_1y_2 ) }\\&=\sqrt{2-2(x_1x_2+y_1y_2)}\end{align}\] Analogously, you will obtain the following: \[|b-c|=\sqrt{2-2(x_2x_3+y_2y_3)} \]\[ |c-a|=\sqrt{2-2(x_3x_1-y_3y_1})\] Now if you substitute every single x and y with the formulas in \((*)\): Just focusing on the term: \(x_1x_2\) \[ \begin{align}x_1x_2&= [-(x_2+x_3)][-(x_1+x_3)]\\&=(x_2+x_3)(x_1+x_3) \\&= x_2x_1+x_2x_3+x_3x_1+x_3^2 \end{align}\] Thus, \(y_1y_2= y_2y_1+y_2y_3+y_3y_1+y_3^2\) In a similar way, you will find: \(x_2x_3=x_1^2+x_1x_2+x_3x_1+x_3x_2\) \(x_3x_1=x_1x_2+x_1x_3+x_2^2+x_2x_3\) Thus: \[ |a-b|=\sqrt{2-2[(x_2x_1+x_2x_3+x_3x_1+x_3^2) +(y_2y_1+y_2y_3+y_3y_1+y_3^2)]}\] \[ |b-c|=\sqrt{2-2[(x_1^2+x_1x_2+x_3x_1+x_3x_2) +(y_1^2+y_1y_2+y_3y_1+y_3y_2)]}\] \[ |c-a|=\sqrt{2-2[(x_1x_2+x_1x_3+x_2^2+x_2x_3) +(y_1y_2+y_1y_3+y_2^2+y_2y_3)]} \]

OpenStudy (kirbykirby):

if you combine \(x_1^2+y_1^2=1\) \(x_2^2+y_2^2=1\) \(x_3^2+y_3^2=1\) Then all the expression are the same

OpenStudy (kirbykirby):

Sorry I made a typo here: \[|c-a|=\sqrt{2-2(x_3x_1+y_3y_1})\]

OpenStudy (zarkon):

it is much simpler to use the hint provided (let a=1) rotations of the circle (the vectors) does not change their length nor the length of their differences

OpenStudy (kirbykirby):

It's still valid though what I wrote, I hope?

OpenStudy (zarkon):

probably ...i didn't read through it ;) I was working on other stuff

myininaya (myininaya):

Wouldn't get a counterexample if we let (a,b,c)=(1,1,-1)?

myininaya (myininaya):

oops the sum is 0 not 1

myininaya (myininaya):

my bad

OpenStudy (zarkon):

I get the lengths of the differences to be \(\sqrt{3}\)

OpenStudy (zarkon):

you can do the entire problem by using geometry on the unit circle

OpenStudy (anonymous):

thank yall!